OpenStudy (annie12m):

If you get it right I medal and fan (ONLY TO THE FIRST PERSON) The product of 3 and a number is 12 more than the sum of the number and 26. What is the number?

OpenStudy (annie12m):

@neonumbrella5115 @Ray1998 @safyousuff97 @Vickyisthesmartone123

OpenStudy (daisyduck04):

easy

OpenStudy (annie12m):

then what is it?

OpenStudy (annie12m):

how do you know?

OpenStudy (daisyduck04):

is the answer choice there

OpenStudy (annie12m):

nope

OpenStudy (daisyduck04):

ok then

OpenStudy (annie12m):

how did you figure it out

OpenStudy (annie12m):

how did you figure it out

OpenStudy (daisyduck04):

ok so what r the answer choices

OpenStudy (photon336):

The product of 3 and a number is 12 more than the sum of the number and 26. let's translate as we go. the product of 3 and n means 3 times n 3*n is means equals 3n = 12 more than the sum of a number and 26 sum of a number n and 26 is (n+26) more than means addition so 12+(n+26) so that leads to our final equation shown below: 3n = 12+(n+26) now we do some basic algebra 38 = 2n 19 = n let's see if this satisfies this. 3(n) should be 12 more than (n+26) 3(19) = 57 right? so let's look at the other side 19+26 = 45 so on one side we have 45 and on the other we have 57 so 57 should be 12 more than 45, if it isn't then our answer is wrong 45+12 = 57 we find that 45+12 does give us 57, so this leads me to believe that the number is 19

OpenStudy (photon336):

@annie12m also this should be in math BTW